# ACEBench: Who Wins the Match Point in Tool Usage?

### v6

@@ 1 Introduction
Large Language Models (LLMs), such as GPT-4 Achiam et al. (2023), have demonstrated exceptional performance across numerous natural language processing tasks Naveed et al. (2023); Qu et al. (2025); Mialon et al. (2023). Studies have shown that incorporating tools can significantly expand LLM capabilities, particularly in specialized domains such as mathematics Das et al. (2024); Bulusu et al. (2024); Gou et al. (2023); Veerendranath et al. (2024), programming Xu et al. (2024), and reasoning Chen et al. (2022); Shao et al. (2022); Surís et al. (2023); Yang et al. (2023). On one hand, integrating tools into LLMs can enhance capabilities in multiple domains, for example, ToolTransformer Schick et al. (2023) enhances the ability of LLMs to solve complex problems by utilizing tools. On the other hand, adopting a tool usage paradigm can improve the robustness of the response and the transparency of the generation, thus increasing the explainability and trust of usersusers Schick et al. (2023), as well as improving the system’s adaptability. As this field continues to evolve, it is essential to comprehensively evaluate all aspects of tool usage, particularly in complex scenarios.
@@
MetaTool Huang et al. (2023)
API-Bank Li et al. (2023)
Stable ToolBench Guo et al. (2024)
BFCL Yan et al. (2024)
$\tau$-Bench Yao et al. (2024)
HammerBench Wang et al. (2024a)
@@
While several studies have focused on evaluating tool usage Yan et al. (2024); Guo et al. (2024); Wang et al. (2024a); Qin et al. (2023); Wang et al. (2024b); Zhuang et al. (2023); Lu et al. (2024), there are still some shortcomings in the existing tool-use benchmarks. Firstly, existing benchmarks lack multi-turn dialogue evaluation in real-world scenarios. For example, the multi-turn dialogues in BFCL Yan et al. (2024) and HammerBench  Wang et al. (2024a) are composed of predefined fixed content combinations. Secondly, current tool-use benchmarks Qin et al. (2023); Guo et al. (2024); Huang et al. (2023); Li et al. (2023) lack fine-grained evaluation and personalized data assessment. Additionally, existing benchmarks Qin et al. (2023); Guo et al. (2024); Wang et al. (2024b) ignore the assessment of special cases, or the evaluation methods are simplistic Yan et al. (2024), as user instructions in real life are not always perfectWang et al. (2024c). The model’s ability to recognize and handle these issues is also crucial for evaluation. Lastly, evaluation costs are high  Qin et al. (2023); Guo et al. (2024), as many studies rely on advanced large models for evaluation.
@@ 2 Related Works
Stable ToolBench Guo et al. (2024) addresses the issue of unstable external APIs by employing a virtual API server, but its dependence on large models for evaluation results in high costs and scalability challenges. BFCL Yan et al. (2024) introduces a benchmark for tool use in multi-turn dialogue scenarios. Yet, it assembles dialogues from fixed content, failing to capture the dynamic and adaptive nature of real-world interactions. Similarly, $\tau$-Bench Yao et al. (2024) evaluates language agents’ ability to engage with human users while adhering to domain-specific rules. Still, its narrow focus on just two scenarios limits its generalizability across diverse tasks. HammerBench Wang et al. (2024a) improves upon this by incorporating datasets derived from popular mobile applications and merging dialogues to simulate typical question-answer trajectories. However, like BFCL, its multi-turn dialogues are simplistic concatenations of pre-defined content, which do not reflect the complexities of real-world conversational dynamics. In addition, some benchmarks Qin et al. (2023); Guo et al. (2024) rely on large language models (LLMs) for result evaluation, leading to high costs and unstable operations.
@@ 3.1.1 Data Construction
(1) API Synthesis. We use real APIs from various real-world scenarios as reference during construction to enhance authenticity. To ensure the stability of the data, we use synthetic APIs to construct the evaluation dataset, referencing real-world APIs as a guide. We employ a self-evolution approach by building a hierarchical API context tree to ensure the generated APIs cover a wide range of domains and functionalities Liu et al. (2024b). Initially, we extract relevant information from technical documents to guide the API generation. As the process progresses, the context tree is gradually expanded, ultimately ensuring the depth and breadth of the generated APIs.
(2) Dialogue Construction. We use two different dialogue generation pipelines built on the constructed API pool from which three to six candidate APIs are selected for each evaluation instance. For most cases, APIs are chosen randomly. However, for instances requiring specific functionality (e.g., similar APIs or multi-turn scenarios), advanced methods, including graph-based sampling Wang et al. (2024d), are used. Simple cases or those with predefined functionality use a template-based generation, where a single generator produces dialogues to ensure consistency. We employ a multi-agent dialogue pipeline for more complex scenarios, where three agents (user, assistant, and tool) to simulate real-world interactions. Both pipelines are supported by carefully hand-crafted examples to ensure comprehensive coverage and diversity.
@@ 4 Experiments
Experimental Setup. In our evaluation, we examine seven closed-source LLMs, including the GPT-4 series Achiam et al. (2023), Qwen-Max Yang et al. (2024), Gemini-1.5-Pro Team et al. (2024), Claude-3.5-Sonnet Anthropic (2024), and Doubao-Pro-32K ByteDance (2025). Additionally, a wide range of open-source language models are assessed, such as the Qwen2.5 series Yang et al. (2024), Llama3 series Dubey et al. (2024), Phi-3-Mini Abdin et al. (2024), Deepseek-V3Liu et al. (2024a), and DeepSeek-Coder-V2 Zhu et al. (2024). Furthermore, four tool-learning-enhanced models were evaluated: Hammer2.1-3B, Hammer2.1-7B Lin et al. (2024), xLAM-7B-r Liu et al. (2024c), and Watt-Tool-8B Watt-AI (2024).
@@ 4.1 Main results and analysis
Loss of Generalization in Fine-Tuned Models. As shown in Figure 3, models fine-tuned on specific datasets, such as Watt-Tool-8B Watt-AI (2024), xLAM-7B Liu et al. (2024c), and Hammer2.1-7BLin et al. (2024), exhibit a significant decline in performance on the Special dataset. This decline can primarily be attributed to the fact that while fine-tuning enhances a model’s performance on specialized tasks, it can also lead to a loss of generalization, making the model less effective on new or broader instruction-following tasks.This phenomenon highlights the importance of balancing task-specific performance and generalization capability during model optimization.
@@ A.1 Data Categories Description
The Special Data refers to situations where the model is unable to resolve the problem posed in the instruction using the candidate functionsWang et al. (2024c). It is categorized into the following categories: Incomplete, Error, and Irrelevant.
@@ B.2 Special Data
{mdframed}
Please refer to the example and continue the dialogue based on the given tool definition and the beginning of the conversation. The requirements are as follows:
1 The user’s request is meaningful, requiring the use of one tool, and the tool will be called once.
2 Calling the tool to fulfill the user’s request still lacks one or more required parameters.
Here is an example for reference
@@
Here is the dialogue continuation you need to write:
<Tool Definition>
{tool_definition}
@@
{mdframed}
Please refer to the example and continue the dialogue based on the given tool definition and the beginning of the conversation. The requirements are as follows:
1 The user’s request is specific and meaningful, requiring one tool, and the tool should be called once.
2 The user’s first question contains incorrect parameters (i.e., it does not conform to the function definition’s pattern or format requirements).
Here is an example for reference
<tool_definition>
@@
Here is the dialogue continuation you need to write:
<Tool Definition>
{tool_definition}
@@ Appendix E Formula for Overall Accuracy
All $\text{All Acc}=A\cdot\text{Acc}_{\text{Normal}}+B\cdot\text{Acc}_{\text{Special}}+C\cdot\text{Acc}_{\text{Agent}}$
@@ G.2 Error Examples of Special Data
{mdframed}
As a user, your role is to interact with an agent. However, during the interaction, you need to follow these guidelines:
1 Break down your inquiries and only raise one question per exchange to simulate a real user’s messages.
2 Provide all the necessary information for the current step. For instance, when setting a reminder, you must give details such as the reminder’s description, title, and time.
3 When asked if you require further assistance, ensure that the main tasks in the instruction have been completed. If not, continue to present the next step to the agent.
4 When the agent asks which message needs to be deleted, proceed with the deletion as specified in the instructions. You cannot offer proactive help to the agent; respond to the agent’s questions according to the instructions, and do not invent any information that you do not know.
5 Once all tasks are complete, generate a ’finish conversation’ message as a standalone line to end the discussion.
Question: question

### v8

@@ 1 Introduction
Large Language Models (LLMs), such as GPT-4 gpt4, have demonstrated exceptional performance across numerous natural language processing tasks llmsurvey; toolsurvey; survey2. Studies have shown that incorporating tools can significantly expand LLM capabilities, particularly in specialized domains such as mathematics mathsensei; math_bulusu2024mathviz; math_gou2023tora; math_veerendranath2024calc, programming xu2024core, and reasoning reason_chen2022program; reason_shao2022chaining; reason_suris2023vipergpt; reason_yang2023mm. On one hand, integrating tools into LLMs can enhance capabilities in multiple domains, for example, ToolTransformer tooltransformer enhances the ability of LLMs to solve complex problems by utilizing tools. On the other hand, adopting a tool usage paradigm can improve the robustness of the response and the transparency of the generation, thus increasing the explainability and trust of usersusers tooltransformer, as well as improving the system’s adaptability. As this field continues to evolve, it is essential to comprehensively evaluate all aspects of tool usage, particularly in complex scenarios.
@@
MetaTool metatool
API-Bank apibank
Stable ToolBench stabletoolbench
BFCL bfcl
$\tau$-Bench taubench
HammerBench hammerbench
@@
While several studies have focused on evaluating tool usage bfcl; stabletoolbench; hammerbench; toolllm; mtubench; toolqa; toolsandbox, there are still some shortcomings in the existing tool-use benchmarks. Firstly, existing benchmarks lack multi-turn dialogue evaluation in real-world scenarios. For example, the multi-turn dialogues in BFCL bfcl and HammerBench  hammerbench are composed of predefined fixed content combinations. Secondly, current tool-use benchmarks toolllm; stabletoolbench; metatool; apibank lack fine-grained evaluation and personalized data assessment. Additionally, existing benchmarks toolllm; stabletoolbench; mtubench ignore the assessment of special cases, or the evaluation methods are simplistic bfcl, as user instructions in real life are not always perfectnosiy_toolbench. The model’s ability to recognize and handle these issues is also crucial for evaluation. Lastly, evaluation costs are high  toolllm; stabletoolbench, as many studies rely on advanced large models for evaluation.
@@ 2 Related Works
Stable ToolBench stabletoolbench addresses the issue of unstable external APIs by employing a virtual API server, but its dependence on large models for evaluation results in high costs and scalability challenges. BFCL bfcl introduces a benchmark for tool use in multi-turn dialogue scenarios. Yet, it assembles dialogues from fixed content, failing to capture the dynamic and adaptive nature of real-world interactions. Similarly, $\tau$-Bench taubench evaluates language agents’ ability to engage with human users while adhering to domain-specific rules. Still, its narrow focus on just two scenarios limits its generalizability across diverse tasks. HammerBench hammerbench improves upon this by incorporating datasets derived from popular mobile applications and merging dialogues to simulate typical question-answer trajectories. However, like BFCL, its multi-turn dialogues are simplistic concatenations of pre-defined content, which do not reflect the complexities of real-world conversational dynamics. In addition, some benchmarks toolllm; stabletoolbench rely on large language models (LLMs) for result evaluation, leading to high costs and unstable operations.
@@ 3.1.1 Data Construction
(1) API Synthesis. We use real APIs from various real-world scenarios as reference during construction to enhance authenticity. To ensure the stability of the data, we use synthetic APIs to construct the evaluation dataset, referencing real-world APIs as a guide. We employ a self-evolution approach by building a hierarchical API context tree to ensure the generated APIs cover a wide range of domains and functionalities toolace. Initially, we extract relevant information from technical documents to guide the API generation. As the process progresses, the context tree is gradually expanded, ultimately ensuring the depth and breadth of the generated APIs.
(2) Dialogue Construction. We use two different dialogue generation pipelines built on the constructed API pool from which three to six candidate APIs are selected for each evaluation instance. For most cases, APIs are chosen randomly. However, for instances requiring specific functionality (e.g., similar APIs or multi-turn scenarios), advanced methods, including graph-based sampling toolflow, are used. Simple cases or those with predefined functionality use a template-based generation, where a single generator produces dialogues to ensure consistency. We employ a multi-agent dialogue pipeline for more complex scenarios, where three agents (user, assistant, and tool) to simulate real-world interactions. Both pipelines are supported by carefully hand-crafted examples to ensure comprehensive coverage and diversity.
@@ 4 Experiments
Experimental Setup. In our evaluation, we examine seven closed-source LLMs, including the GPT-4 series gpt4, Qwen-Max qwen2.5, Gemini-1.5-Pro gemini1.5, Claude-3.5-Sonnet claude-3.5, and Doubao-Pro-32K doubao-pro-32k. Additionally, a wide range of open-source language models are assessed, such as the Qwen2.5 series qwen2.5, Llama3 series llama3, Phi-3-Mini phi3, Deepseek-V3deepseek-v3, and DeepSeek-Coder-V2 DeepSeek-Coder-V2. Furthermore, four tool-learning-enhanced models were evaluated: Hammer2.1-3B, Hammer2.1-7B hammer, xLAM-7B-r xlam, and Watt-Tool-8B watt-tool.
We additionally report the results of Kimi-K2 kimik2 and LoopTool-8B looptool on the English dataset.
@@ 4.1 Main results and analysis
Loss of Generalization in Fine-Tuned Models. As shown in Figure 3, models fine-tuned on specific datasets, such as Watt-Tool-8B watt-tool, xLAM-7B xlam, and Hammer2.1-7Bhammer, exhibit a significant decline in performance on the Special dataset. This decline can primarily be attributed to the fact that while fine-tuning enhances a model’s performance on specialized tasks, it can also lead to a loss of generalization, making the model less effective on new or broader instruction-following tasks.This phenomenon highlights the importance of balancing task-specific performance and generalization capability during model optimization.
@@ A.1 Data Categories Description
The Special Data refers to situations where the model is unable to resolve the problem posed in the instruction using the candidate functionsnosiy_toolbench. It is categorized into the following categories: Incomplete, Error, and Irrelevant.
@@ B.2 Special Data
Please refer to the example and continue the dialogue based on the given tool definition and the beginning of the conversation. The requirements are as follows: 1 The user’s request is meaningful, requiring the use of one tool, and the tool will be called once. 2 Calling the tool to fulfill the user’s request still lacks one or more required parameters. …… Here is an example for reference
@@
Here is the dialogue continuation you need to write: <Tool Definition> {tool_definition} ……
@@
Please refer to the example and continue the dialogue based on the given tool definition and the beginning of the conversation. The requirements are as follows: 1 The user’s request is specific and meaningful, requiring one tool, and the tool should be called once. 2 The user’s first question contains incorrect parameters (i.e., it does not conform to the function definition’s pattern or format requirements). …… Here is an example for reference <tool_definition>
@@
Here is the dialogue continuation you need to write: <Tool Definition> {tool_definition} ……
@@ Appendix E Formula for Overall Accuracy
$\text{All Acc}=A\cdot\text{Acc}_{\text{Normal}}+B\cdot\text{Acc}_{\text{Special}}+C\cdot\text{Acc}_{\text{Agent}}$
@@ G.2 Error Examples of Special Data
As a user, your role is to interact with an agent. However, during the interaction, you need to follow these guidelines: 1 Break down your inquiries and only raise one question per exchange to simulate a real user’s messages. 2 Provide all the necessary information for the current step. For instance, when setting a reminder, you must give details such as the reminder’s description, title, and time. 3 When asked if you require further assistance, ensure that the main tasks in the instruction have been completed. If not, continue to present the next step to the agent. 4 When the agent asks which message needs to be deleted, proceed with the deletion as specified in the instructions. You cannot offer proactive help to the agent; respond to the agent’s questions according to the instructions, and do not invent any information that you do not know. 5 Once all tasks are complete, generate a ’finish conversation’ message as a standalone line to end the discussion. Question: question